OpenStudy (anonymous):

\[evaluated the interval \int\limits_{}^{} x/(x-1)^{3}dx\]

OpenStudy (anonymous):

\[\frac{x}{(x-1)^3}=\frac{1}{(x- 1)^2}+\frac{1}{(x-1)^3} \] By Partial fraction, then it is easy.

OpenStudy (anonymous):

\[ \int \frac{1}{(x-1)^3} \, dx = -\frac{1}{2 (x-1)^2} +C \]

OpenStudy (anonymous):

\[ \int \frac{1}{(x-1)^2} \, dx =-\frac{1}{x-1} + C \]

OpenStudy (anonymous):

Combining the tw and simplifying, you get the answer as \[ \frac{1-2 x}{2 (x-1)^2} + C \]

OpenStudy (anonymous):

That is what we combined to get the final answer \[-\frac{1}{x-1}-\frac{1}{2 (x-1)^2} \]

OpenStudy (anonymous):

Use u-substitution take \[u=x-1\] then \[du=dx\] This simplifies the the denominator. Now substitute u in....u=x-1 and x=u+1 Now you have: \[\int\limits_{}^{} ((u+1)/u^{3})du\] (I can't figure out how to show my division as a fraction on this site yet, sorry) Now break up your fraction into two different ones \[\int\limits_{}^{} ((u/u^{3})+(1/u ^{3}))du\] Simplify: \[\int\limits_{}^{} ((1/u^{2})+(1/u ^{3}))du\] Now you can take the integral of the two fractions individually. \[\int\limits_{}^{} (1/u^2)du=-1/u+c_{1}\] \[\int\limits_{}^{} (1/u^3)du=-1/(2u^2)+c_{2}\] Don't forget to add the two fractions: \[-1/u+-1/(2u^2)+c\] Simplify (common denominators): \[(-2u-1)/(2u^2)+c\] Substitute x-1 back in for u \[(-2(x-1)-1)/(2(x-1)^2)\] \[(-2x+2-1)/(2(x-1)^2)\] \[(-2x+1)/(2(x-1)^2)\] (Again, sorry about the notation. How do you represent division as a fraction using this site?)

OpenStudy (zarkon):

to do fractions use \frac{a}{b} \[\frac{a}{b}\]